In a binomial experiment the probability of success for females is 0.7. What is the probability of three successes in seven trials for females? a.0.082 b.0.097 c.0.227 d.0.185
Answer #1

Solution

Given that ,

p = 0.7

q = 1 - p = 1 - 0.7=0.3

n = 7

Using binomial probability formula ,

P(X = x) = (n C x) * p x * (1 - p)n - x

P(X = 3) = (7 C 3) * 0.7 3 * (0.3)4

=0.0972

probability=0.097